# Improving Model Evaluation using SMART Filtering of Benchmark Datasets

### v1

@@ 2 Related Works
Other works have explored related approaches to improve evaluation Bras et al. (2020). Varshney et al. (2022) tested models only on difficult examples of a dataset, showing that evaluating on just 5% examples can achieve a very high correlation with the full dataset. However, their approach requires retraining the models on various subsets of training data and uses humans to verify the filtered dataset. In the current NLP paradigm, access to training data is often unavailable for most models Longpre et al. (2024) and even when accessible, it is very expensive to retrain models on those datasets. Contrasting with Varshney et al. (2022), our approach emphasizes computational efficiency and identifies a high-quality subset that correlates well with the original dataset, without requiring access to training data or expensive human verification.
@@ 3.2 Filtering Easy Examples
In this step, we focus on identifying and removing the easy examples from the dataset Gururangan et al. (2018); Rodriguez et al. (2021). We define easy examples based on the agreement between top-performing open-source models from Open LLM leaderboard (HuggingFace, [2024]). Specifically, we consider an example to be easy if there is unanimous agreement among all top-performing models, with each model answering the example correctly with high confidence (greater than $0.8$). This approach finds examples that are consistently easy across different model architectures and models trained on different datasets. By removing these examples, we also reduce the computational cost of running inference on the resulting dataset.
@@ 3.3 Filtering Data Contaminated Examples
In this step, we identify and remove examples that are likely to be data contaminated, i.e. present in training data of models Deng et al. (2023); Elazar et al. (2024); Magar and Schwartz (2022). This is important because evaluating models on examples they have already seen during training can lead to inflated numbers, giving a false sense of higher model performance. To identify data contaminated examples, we follow an approach inspired by Balepur et al. (2024) and Balepur and Rudinger (2024) as shown in Figure 2. Specifically, we modify the prompt to the model by removing the question text and presenting only the answer choices. This approach challenges the model to select the correct answer in an artificial setting without the context of the question itself.
In this step, we again take agreement between top-performing models. Our hypothesis is that if all the top n open-source models can predict the correct answer with a high confidence (greater than $0.8$) for answer-only setting, then it is highly likely that the example has been seen during the training by these models and is therefore data contaminated.
@@ 4 Results: Efficiency & Informativeness
In this section, we present the results of applying SMART filtering on popular datasets. Our methodology is generic and can be applied to various dataset types. We demonstrate its efficacy using three well-established multiple-choice question answering datasets: ARC Clark et al. (2018), MMLU Hendrycks et al. (2021a) and CommonsenseQA Talmor et al. (2019). We release our code and provide SMART filtering version of each dataset. [^1] We will give link to our codebase upon paper acceptance.
@@
The impact of SMART filtering varies significantly across datasets. In ARC, 64.4% of the examples were classified to be ‘easy’, compared to only 27.6% in CommonsenseQA. In MMLU, a significant 4.37% of the dataset was found to be data contaminated, while CommonsenseQA had none. Additionally, 12.57% of ARC and 4.56% of MMLU examples were identified as similar. These findings highlight that different datasets might suffer from different types of low-quality examples. For MMLU, a widely accepted dataset for measuring LLM performance, we were able to filter out 43% of the dataset while maintaining almost the same model ranking. This suggests that MMLU might not be saturated in the end, and there is still room for further hill-climbing for current models.
While relative model ranking is maintained, SMART filtering leads to a notable decrease in inter-model agreement across all datasets, as measured by the percentage of samples yielding identical predictions among models (Figure 4). This reduction in model agreement is beneficial, as it enables better differentiation between models based on their capabilities, and provides more headroom for improvement by decreasing their correlation.
@@
4.1 Category Wise Results
Our analysis reveals significant variation in the percentage of examples removed by SMART filtering across different categories in MMLU. Over 60% of examples were removed in 9 categories, while less than 20% were removed in $13$ categories. Notably, ‘high school government and politics’, ‘high school psychology’, ‘marketing’, and ‘sociology’ saw a reduction in size by 73%, 67%, 63%, and 62% respectively. In contrast, ‘abstract algebra’, ‘global facts’, ‘high school physics’, and ‘professional law’ were much less affected, with only 4%, 4%, 5%, and 12.6% of examples removed, respectively. A detailed breakdown is shown in Appendix A.4.
State-of-the-art models often report accuracy on the ARC-Challenge category of ARC, as it is presumed to be more difficult Meta AI (2024); Google (2024). Suprisingly, our analysis shows no significant difference between ARC-Easy and ARC-Challenge after SMART filtering. Both categories were significantly reduced, with 73% of ARC-Easy and 60% of ARC-Challenge examples removed. A high proportion, 55.1% of ARC-Challenge were deemed easy by all models.
4.2 Correlation in Model Rankings
@@ 5 Robustness of our Approach
In this section, we show that SMART filtering is robust to altering settings, such as (i) the number of models used for identifying easy and data contaminated examples and (ii) choice of embedding model used to find similar examples.
@@
5.1 Ablation: Number of Models for Easy and Data Contaminated Examples
In the SMART filtering methodology, we used 7 models for identifying easy and data contaminated examples. Our selection was based on selecting the top 7 opensource models from different organizations. To assess the robustness of our approach, we conducted an ablation study varying the number of models used. We randomly selected subsets of 4, 5, and 6 models from the original set of 7. Specifically, we run our methodology on 10 random subsets of 4 models from a total of 35 subsets ($\binom{7}{4}=35$), 10 random subsets of 5 models from a total of 21 subsets ($\binom{7}{5}=21$) and all 7 subsets of 6 models ($\binom{7}{6}=7$). The results, presented in Table 4, show that the percentage of examples filtered remains relatively stable across different model combinations for all three datasets. This shows that our methodology is largely insensitive to the number of models used, and can be effectively applied using any n top performing open-source models, again opening up possibilities for efficiency gains.
@@ 6 Discussion
Our approach removes the easy, data contaminated, and similar examples, thus improving the quality of the resulting dataset. Our methodology can be applied at various stages, including before or after dataset release. SMART filtering can also be iteratively applied to benchmark datasets to maintain quality and adapt to future model releases.
@@
Recent studies have explored testing models on dataset subsets while maintaining high correlation with original dataset performance Varshney et al. (2022). Our approach achieves substantial dataset size reductions, up to 68.9% for ARC, while preserving similar model rankings as shown in Tables 1 and 3. These results are linearly correlated, 68.9% decreases in computation time and evaluation costs for testing new models on the ARC dataset. Our methodology thus offers significant time and cost savings for future benchmark evaluations, aligning with the growing need for efficient model assessment techniques.
@@ A.7 Detailed Model Results
Table 8, 9, and 10 shows accuracy for all tested models on ARC, MMLU, and CommonsenseQA dataset and their SMART filtered versions.

### v2

@@ 2 Related Works
Other works have explored related approaches to improve evaluation Vivek et al. (2024); Bras et al. (2020); Gupta et al. (2023). Varshney et al. (2022) tested models only on difficult examples of a dataset, showing that evaluating on just 5% examples can achieve a very high correlation with the full dataset. However, their approach requires retraining the models on various subsets of training data and uses humans to verify the filtered dataset. In the current NLP paradigm, access to training data is often unavailable for most models Longpre et al. (2024) and even when accessible, it is very expensive to retrain models on those datasets. Contrasting with Varshney et al. (2022), our approach emphasizes computational efficiency and identifies a high-quality subset that correlates well with the original dataset, without requiring access to training data or expensive human verification.
@@ 3.2 Filtering Easy Examples
In this step, we focus on identifying and removing the easy examples from the dataset Gururangan et al. (2018); Rodriguez et al. (2021a). We define easy examples based on the agreement between top-performing open-source models from Open LLM leaderboard (HuggingFace, [2024]). Specifically, we consider an example to be easy if there is unanimous agreement among all top-performing models, with each model answering the example correctly with high confidence (greater than $0.8$). This approach finds examples that are consistently easy across different model architectures and models trained on different datasets. By removing these examples, we also reduce the computational cost of running inference on the resulting dataset.
@@ 3.3 Filtering Data Contaminated Examples
In this step, we identify and remove examples that are likely to be data contaminated, i.e. present in training data of models Deng et al. (2023); Elazar et al. (2024); Magar and Schwartz (2022). This is important because evaluating models on examples they have already seen during training can lead to inflated numbers, giving a false sense of higher model performance. Detecting data contamination is inherently challenging as also highlighted in recent studies Duan et al. (2024); Singh et al. (2024), and many promising approaches require access to training data Jiang et al. (2024). To identify data contaminated examples, we follow an approach inspired by Balepur et al. (2024) and Balepur and Rudinger (2024) as shown in Figure 2. In this approach, we do not need access to training data. Specifically, we modify the prompt to the model by removing the question text and presenting only the answer choices. This approach challenges the model to select the correct answer in an artificial setting without the context of the question itself.
@@
In this step, we again take agreement between top-performing models. We employ a stringent criterion: an example is deemed contaminated only if all top n open-source models answer it correctly without any context and with high confidence (greater than $0.8$). This conservative approach prioritizes precision over recall to ensure that the examples removed are truly contaminated. We recognize that this strict criterion may underestimate the number of contaminated examples. As the research in detecting data contamination improves, we plan to incorporate more sophisticated techniques that can detect contamination more effectively while balancing precision and recall.
@@ 4 Results: Efficiency & Informativeness
In this section, we present the results of applying SMART filtering on popular datasets, focusing on three well-established multiple-choice question answering datasets: ARC Clark et al. (2018), MMLU Hendrycks et al. (2021a) and CommonsenseQA Talmor et al. (2019). While we show results on MCQA datasets, our methodology is generic and can be applied to various dataset types. Given that all other classification tasks can be, and often are, recast into MCQA format, it strikes us as likely that SMART-filtering would work just as well for classical NLU tasks (like NLI, coreference resolution, sentiment analysis) without needing novel strategies. Although these tasks have conceptual differences, with respect to the format, these variations boil down to having different numbers of answer options. Going into specifics of why we selected these datasets, we selected MMLU and ARC datasets because of their widespread adoption in benchmarking by state-of-the-art models Meta AI (2024); Google (2024) and their coverage of diverse topics and difficulties. These two datasets are 4-choice MCQA, so to broaden the scope, we included CommonsenseQA, which offers 5-choice MCQA. We think our inclusion of different numbers of MCQA answer options between 4 and 5 gives some indication that the methodology will be adaptable to smaller numbers of answer options (2 or 3) too. We release our code and provide SMART filtering version of each dataset111 Link to our codebase: https://github.com/facebookresearch/ResponsibleNLP/tree/main/SMART-Filtering..
@@
The impact of SMART filtering varies significantly across datasets. In ARC, 64.4% of the examples were classified to be ‘easy’, compared to only 27.6% in CommonsenseQA. In MMLU, a significant 4.37% of the dataset was found to be data contaminated, while CommonsenseQA had none. Additionally, 12.57% of ARC and 4.56% of MMLU examples were identified as similar. These findings highlight that different datasets might suffer from different types of low-quality examples. For MMLU, a widely accepted dataset for measuring LLM performance, we were able to filter out 43% of the dataset while maintaining almost the same model ranking. This suggests that MMLU might not be saturated yet, and there is still room for further hill-climbing for current models.
While relative model ranking is maintained, SMART filtering leads to a notable decrease in inter-model agreement across all datasets, as measured by the percentage of samples yielding identical predictions among models (Figure 4). This reduction in model agreement is beneficial, as it enables better differentiation among models based on their capabilities, and provides more headroom for improvement by decreasing their correlation.
@@
4.1 Category-Wise Results
Our analysis reveals significant variation in the percentage of examples removed by SMART filtering across different categories in MMLU. Some categories appear to be more challenging than others. Over 60% of examples were removed in 9 categories. Notably, ‘high school government and politics’, ‘high school psychology’, ‘marketing’, and ‘sociology’ saw a reduction in size by 73%, 67%, 63%, and 62% respectively. A qualitative look into examples of these categories indicates they contain questions with well-known answers, making them easier for models and more susceptible to be present during training and possible data contamination, while less than 20% were removed in $13$ categories. ‘Abstract algebra’, ‘global facts’, ‘high school physics’, and ‘professional law’ have only 4%, 4%, 5%, and 12.6% of examples removed, respectively. These subjects require complex reasoning and problem-solving, posing greater challenges for models. A detailed breakdown is shown in Appendix A.4.
State-of-the-art models often report accuracy on the ARC-Challenge category of ARC, as it is presumed to be more difficult Meta AI (2024); Google (2024). Surprisingly, our analysis shows no significant difference between ARC-Easy and ARC-Challenge after SMART filtering. Both categories were significantly reduced, with 73% of ARC-Easy and 60% of ARC-Challenge examples removed. A high proportion, 55.1% of ARC-Challenge were deemed easy by all models.
4.2 Correlation in Model Rankings
@@ 4.3 Correlation between ChatBot Arena scores and Model Accuracy
Dataset Name Original SMART Random IRE
ARC 0.783 0.845 0.774 0.784
MMLU 0.764 0.776 0.767 0.766
CommonsenseQA 0.666 0.660 0.659 0.658
Table 4: Comparison of performance metrics across different methods. Random signifies random baseline as mentioned in Section 4.4.
@@
4.4 Comparison with Other Approaches
In this section we compare SMART filtering with other similar approaches and a random baseline. Rodriguez et al. (2021b) introduce a method that filters examples based on difficulty and discriminability, aiming to make leaderboards more informative. For this baseline, we used 25th percentile thresholds for difficulty and discriminability criteria to establish this baseline. We call their method IRE. We also provide analysis with a random baseline where we created a random baseline by reducing each dataset to the same size as our SMART-filtered datasets (reducing dataset size randomly to 69%, 43% and 34% for ARC, MMLU and CommonsenseQA).
Dataset Name % of Examples Filtered
SMART IRE
ARC 68.9 28.1
MMLU 43.0 31.1
CommonsenseQA 34.3 27.8
Table 5: Percentage of examples filtered by SMART vs. IRE for each dataset.
As seen in Table 4, SMART Filtering consistently achieves a higher correlation with human preferences than both the random baseline and IRE. Additionally, SMART Filtering consistently filters out a larger portion of the datasets compared to IRE, leading to greater efficiency in evaluation (c.f. Table 5).
@@ 5 Robustness of our Approach
In this section, we show that SMART filtering is robust to altering settings, such as (i) the number of models used for identifying easy and data-contaminated examples and (ii) choice of embedding model used to find similar examples.
@@
5.1 Ablation: Number of Models for Easy and Data-Contaminated Examples
In the SMART filtering methodology, we used 7 models for identifying easy and data-contaminated examples. Our selection was based on selecting the top 7 opensource models from different organizations. To assess the robustness of our approach, we conducted an ablation study varying the number of models used. We randomly selected subsets of 4, 5, and 6 models from the original set of 7. Specifically, we run our methodology on 10 random subsets of 4 models from a total of 35 subsets ($\binom{7}{4}=35$), 10 random subsets of 5 models from a total of 21 subsets ($\binom{7}{5}=21$) and all 7 subsets of 6 models ($\binom{7}{6}=7$). The results, presented in Table 6, show that the percentage of examples filtered remains relatively stable across different model combinations for all three datasets. This shows that our methodology is largely insensitive to the number of models used, and can be effectively applied using any n top performing open-source models, again opening up possibilities for efficiency gains.
@@ 5.2 Embeddings for Similar Examples
5.3 Quality of Filtering
To ensure that our method effectively filters out unwanted instances, we manually looked at 100 removed examples from the Easy and Similar examples category. For Similar examples, we found that 93 were genuinely similar, 5% were borderline similar and 2% were not similar. For Filtering Easy Examples, 97 examples could be easily answered by the authors or found via a simple web search. This analysis shows the effectiveness of our approach as most of the filtered out examples are unwanted for comparing model performance.
@@ 6 Discussion
Our approach removes the easy, data contaminated, and similar examples, thus improving the quality of the resulting dataset. Our methodology can be applied at various stages, including before or after dataset release. SMART filtering is designed to be iterative and scalable; we intend to periodically reapply SMART-Filtering to existing benchmarks to identify and filter out examples that become too easy or contaminated for newer models. We suspect that, as models become more capable, what constitutes an easy example could change.
@@
Recent studies have explored testing models on dataset subsets while maintaining high correlation with original dataset performance Varshney et al. (2022). Our approach achieves substantial dataset size reductions, up to 68.9% for ARC, while preserving similar model rankings as shown in Tables 1 and 3. As dataset size correlates with computation time and evaluation costs, the 68.9% decrease in size should be reflected in similar decreases in time and evaluation cost. Our methodology thus offers significant time and cost savings for future benchmark evaluations, aligning with the growing need for efficient model assessment techniques.
@@ A.7 Detailed Model Results
Table 10, 11, and 12 shows accuracy for all tested models on ARC, MMLU, and CommonsenseQA dataset and their SMART filtered versions.